OpenStudy (anonymous):

can you solve for x please 550=((1)/(1/1120)) + 1/x

OpenStudy (mertsj):

\[550=1120+\frac{1}{x}\]

OpenStudy (anonymous):

is that a double fraction? if it is then 550=1120/1 + 1/x 1/x=550- 1120 then u can go from there

OpenStudy (mertsj):

\[550x=1120x+1\]

OpenStudy (anonymous):

You can probably solve it from @Mertsj edit.

OpenStudy (mertsj):

\[-570x=1\]

OpenStudy (mertsj):

\[x=\frac{-1}{570}\]

OpenStudy (anonymous):

550 =1120+1/x 550-1120=1120-1120 +1/x -570 =1/x x =-1/570

OpenStudy (australopithecus):

550=((1)/(1/1120)) + 1/x remember every can be expressed as a fraction divided by 1, so 1 become 1/1 now just flip bottom fraction and multiply (1/1) / (1/1120) + 1/x = 550 = (1120x + 1)/x = 550 1 = 550x - 1120x you can do the rest